# Is Human-Like Text Liked by Humans? Multilingual Human Detection and Preference Against AI

### v1

@@ 5 Human-Like or Liked-by-Human?
We used the prompting strategy to bridge the gap between human-written and machine-generated text, aiming to make machine outputs more human-like. However, do humans favor human-like text? Below, we examine human preferences among four options: (1) human-written text, (2) machine-generated text using the original prompt, (3) machine-generated text using the improved prompt, or (4) none of the above.
@@
Preference Labeling Setup:
We labeled the preferences for three languages: Chinese, Russian, and Arabic. For Chinese, we annotated Zhihu-QA and student essays (300 examples for each), along with 100 responses particularly for Zhihu questions, where emotional and empathetic comforts are highlighted. Five unique annotators participated, identified by nationality-gender-degree. For example, Zh-Male-PhD refers to a Chinese male, who is a PhD student. Similarly, we labeled two datasets for Russian (three annotators) and two datasets for Arabic (two annotators).
Do People Always Prefer Human Text?
The answer is No. Analyzing the preferences of ten annotators across six datasets in Figure 3 and 4, human text was preferred in about half of the cases. Notably, for Russian and Arabic, annotators tended to favor machine text. This is particularly evident for Russian summaries using the improved prompts (green bars) and Arabic summaries using the original prompts (orange bars). For Chinese datasets, including Zhihu QA and student essays, human-written text is generally preferred though there are exceptions. Zh-Male-postdoc exhibits a unique preference distribution that deviates from the rest. Interestingly, for emotion-rich questions (QA-emo), where human responses are expected to be more empathetic and be preferred, three out of four annotators actually prefered the machine-generated responses. The remaining annotator disliked both the human and the machine text in 22% of the cases. The annotator feedback suggested that this was influenced by the presence of mean-spirited responses from Zhihu users, where some human answers expressed personal biases and lacked empathy (see D.12).
@@
Human-written essays exhibit greater coherence and sincerity. In contrast, machine-generated essays often lack cohesion, displaying abrupt transitions and sometimes repeating the title to create superficial continuity. While LLMs can generate related concepts, they struggle to construct a compelling narrative or to seamlessly connect multiple stories under a unified theme. As a result, their outputs often lack depth, relying on abstract concepts and rhetorical flourishes without delivering substantive insights. The text tends to be verbose yet superficial, giving an impression of grandiosity without meaning. Moreover, LLM-generated essays often adopt an instructive tone, resembling answers to questions rather than peer-level discourse. From a literary perspective, a well-written human article may be lively, charming, and playful, or it may present a novel perspective that inspires deep reflection. It might feature a cleverly structured beginning and ending, convey sincerity and heartfelt emotions, or captivate with its elegance and melancholy. With fresh and precise wording, it creates a rich atmosphere, evoking a refined aesthetic and a sense of poetic depth. In contrast, a piece generated by an LLM lacks this literary nuance, leaving little room for contemplation or lingering thoughts. The expression is immediately clear, requiring no further reflection, resembling more of a lecture than an engaging discourse.
Preference Distributions Vary Across Individuals.
The annotators for Russian and Arabic exhibited similar preference distributions, whereas large differences occured for Chinese QA. For example, in the Chinese Zhihu QA, the second annotator selected only seven human-written texts, while the fourth one chose 284 (95%). In the QA-emo, the first and the second annotators prefered human-written responses, whereas the third and the fourth favored machine-generated text. This variance shows the charm of collecting personal preferences and then optimizing models to align with individual philosophies. Our preference annotations can serve as a valuable resource for investigating the relationship between annotator characteristics (e.g., MBTI personality, gender, and age) and preferences. Also the data can guide models to match individual preferences in multilingual contexts.
@@
Human texts are not always preferred. This inspires us to reflect the ultimate goal of building LLMs that are human-like vs. liked-by-human. The goal of being human-like has a single target, i.e., mimicking human behavior, while to be liked-by-humans involves optimization towards billions of local optima, each shaped by individual preferences. Current language models establish a foundation by learning from human data towards being human-like. As they get more advanced, they can further adapt by incorporating personal data, thus transitioning from merely imitating human behavior to aligning with individual user preferences, and moving from human-like to liked-by-human.
@@ 6 Conclusion and Future Work
We conducted a comprehensive study to investigate the capability of humans to detect text generated by SOTA LLMs. Based on 16 datasets spanning 9 languages, 9 domains, and 11 LLMs, 19 expert annotators explored the upper bounds of human detection potential. Accuracy ranged in 50-100%, with an average of 87.6%, revealing that it is not that challenging for experts to identify MGTs. We found that the major gaps between human-written and machine-generated text lie in concreteness, cultural nuances, diversity in length, structure, style, and sentiment, formatting and mixture of languages (with English) in the generations. Prompting by explicitly indicating the distinctions could partially bridge this gap, but cultural nuances and diversity remained a challenge. Yet, humans favored machine-generated text in over half of the cases, particularly when they could not recognize which text was written by a human. In future work, we plan to further explore the subtle relationship between human detection accuracy, preference, and individual characteristics: (i) Are humans more likely to favor human-written text when they can identify it? and (ii) How do an individual’s philosophy, personality, and experience influence their ability to discern between human-written and machine-generated text?
@@
Annotator Bias
All annotators involved in the detection process are advanced NLP researchers, including Msc and PhD students, as well as postdocs, specializing in LLMs. No laymen participated in the annotation process, which ensures that the findings and the conclusions were relevant within the expert domain. It is possible that different results could be obtained if laypersons were involved in the detection tasks. However, given our aim to explore the upper bound of human capability in detecting machine-generated text, the current setting is appropriate. Future work could consider involving a broader range of participants to assess the average detection capability across different population groups.
@@
Linguistic Analysis
Due to page limitations, we were unable to present a detailed linguistic analysis of the text from human-written, original machine-generated, and newly generated outputs. We will perform analysis of vocabulary features, part-of-speech tagging, dependency structures, sentiment analysis, language model perplexity, and other relevant metrics in future work.
@@
Our study engaged human evaluators in distinguishing between human-written and machine-generated texts, while expressing their preference. All annotators gave informed consent, were fully aware of the study’s objectives, and had the right to withdraw at any time. Since this study was designed to be evaluated by experts, all annotators participating in it were advanced NLP researchers (MSc and PhD studnets, as well as postdocs, specializing in LLMs). While this ensures a rigorous assessment of the detection capabilities within an expert domain, we acknowledge that findings may not generalize to laypeople. Future research should include a more diverse range of participants to evaluate human detection performance across various demographic and professional backgrounds.
@@ A.1 Human Detection on MGT
Early studies probing the outputs of less advanced generative models found that human evaluators could detect the differences. For example, Garbacea et al. (2019) reported that individual evaluators obtained 66.61% accuracy when evaluating online product reviews generated by 12 sequential models before 2019, and majority voting of five annotators’ predictions improved to 72.63%. Ippolito et al. (2020) found that trained evaluators were able to detect GPT2-large text 71.4%. However, since the emergent of GPT3, it has become challenging for human evaluators to distinguish between human-written and machine-generated text. Evaluators could guess GPT3-generated stories, news articles and recipes with 50% accuracy (Clark et al., 2021). 203 crowd-soured workers employed from Prolific show 57% average accuracy when discerning news generated by GPT-3.5-turbo and news collected from trusted news organizations, e.g., New York Times, Wall Street Journal, and they obtained 78% on social media comments, both with a wide spread in performance across individuals (Chein et al., 2024). Guo et al. (2023) also showed large differences in detection capability between experts and laymen, where the former is much higher than the later across different domains. Average accuracy of experts is above 90%, but random guess for amateurs, i.e., 48% and 54% for English and Chinese respectively. With expert annotators, academic paper abstracts are harder to detect than the Reddit answers. Detection accuracy ranges from 41 to 94 for Reddit with an average of 77, while 72 for abstracts with individual performance rangining from 60 to 84, revealing that scientific materials are more challenging even for experts who are researchers and frequent users of LLMs Wang et al. (2024b). In addition to binary detection human vs. AI, Wang et al. (2024a) performed 4-class detection to identify which LLM generated a given text, achieving accuracy of 21.2%, below the random guess of 25%, though annotators are all NLP PhD students who are exploring machine-generated text in their research. Dugan et al. (2023) asked annotators to select boundary sentences between human- and machine-authored snippets given a mixed collaborative text by a game tool. They found that players correctly identified the boundary sentence 23.4% of the time (chance being 10%). As shown in Table 1, previous studies primarily focus on English and outputs of GPT-3.5-turbo, non-English languages and more advanced models have been under-investigated. The accuracy of human detection is influenced by languages, generative models, and domains of the given text, as well as the annotators. Large variance is observed among individual evaluators, particularly between experts and laymen. Therefore, we aim to fill this gap by a comprehensive case study exploring whether human can safeguard against AI over 16 dataset spanning nine languages, nine domains and 11 SOTA LLMs in four evaluation setups.
@@ A.3 Detection Accuracy and Preference
What factors may influence individuals’ discerning accuracy? Chein et al. (2024) investigated whether psychological attributes (e.g., fluid intelligence, executaive functionining, empathy) or experience (e.g., smartphone and social media habits) predict the ability to differentiate between human and AI materials. Results show that only fluid intelligence — the capacity to reason and solve novel problems without relying on previously acquired knowledge, was strongly associated with the ability to differentiate human from AI materials. Other three factors were not significantly related to the overall detection accuracy. [^1] Fluid intelligence involves the ability to think abstractly, analyze patterns, make decisions in unfamiliar situations, and adapt to new challenges. Fluid intelligence is typically thought to be influenced by genetic and biological factors, and it tends to peak in early adulthood before gradually declining with age. It’s often contrasted with crystallized intelligence, which is based on accumulated knowledge and experience. They also find that increased exposure to mixed AI and human materials online does not enhance one’s ability to differentiate human from AI materials, but rather, may make AI-generated content appear more human-like. Combine together, language proficiency, cognitive intelligence, and familiarity with AI outputs may affect detection accuracy. Intuitively, a high level of language proficiency will make it easier to spot overly formal, repetitive, or mechanically structured phrasing typical of AI outputs. Native speakers with strong awareness of cultural nuance can better understand cultural and idiomatic expressions, facilitating identify content that lacks human-like subtlety or creativity. Strong analytical skills can help detecting inconsistencies, logical flaws, style patterns that indicate MGT. Individuals with frequent interaction with LLMs may recognize their writing patterns like positive tone bias Guo et al. (2023); Chein et al. (2024). Therefore, to probe the upper bound of human capability of detecting current SOTA LLM generations, all annotators for MGT detection and preference labeling are native speakers for each language, Masters’, who are PhD and postdoc students studying LLMs, with 5 female and 14 male.
Приложение B Datasets
@@ B.1 Arabic
The SANAD news dataset Einea et al. (2019) is an extensive collection of Arabic news articles in Modern Standard Arabic (MSA), designed to support various Arabic NLP tasks. SANAD dataset comprises over $190,000$ articles from three prominent news websites: AlKhaleej, AlArabiya, and Akhbarona, and it is released under the license of CC-BY-4.0. For this study, we generated a total of $500$ news articles using titles from the SANAD news dataset, with $100$ articles for each of five selected news category topics present in SANAD spanning finance, politics, sports, medicine, and technology. The articles were generated using the OpenAI’s GPT-4o-2024-05-13 and the prompt used to generate these articles is shown below:
@@ B.2 Chinese
B.3 English Peer Meta Review
@@ B.4 Hindi News
B.5 Italian DICE News
@@ B.9 Vietnamese
Appendix C Distinguishable Signals
@@ C.2 Chinese
C.3 English Peer Meta Review
@@ C.9 Vietnamese
Appendix D Fill the Gap by Prompting?

### v2

@@ 5 Human-Like or Liked-by-Human?
We used the prompting strategy to bridge the gap between human and machine text, aiming to make machine outputs more human-like. However, do humans favor human-like text? Below, we examine human preferences among four options: (i) human-written text, (ii) machine-generated text using the original prompt, (iii) machine-generated text using the improved prompt, or (vi) none of the above.
@@
Preference Labeling Setup:
To evaluate the impact of human detection accuracy on their preferences, we selected datasets with high accuracy (Chinese QA and essays, Russian news), medium accuracy (Russian and Arabic summaries), and low accuracy (Arabic tweets). For Chinese, we annotated Zhihu-QA and student essays (300 examples for each), along with 100 responses particularly for Zhihu questions, where emotional and empathetic comforts are highlighted. Five unique annotators participated, identified by nationality-gender-degree. For example, Zh-Male-PhD refers to a Chinese male PhD student. We labeled Russian with three annotators and two annotators for Arabic.
Do People Always Prefer Human Text?
The answer is No. Analyzing the preferences of ten annotators across six datasets in Figure 3 and 4, human text was preferred in about half of the cases. Notably, for Russian and Arabic, annotators tended to favor machine text. This is particularly evident for Russian summaries using the improved prompts (green bars) and Arabic summaries using the original prompts (orange bars). Annotators prefer human-written text when they can clearly identify its source, otherwise they tend to favor machine text. For Chinese Zhihu QA and student essays, where detection accuracy approaches 100%, human-written text is generally preferred though there are exceptions. Zh-Male-postdoc exhibits a unique preference distribution that deviates from the rest. Ru-Male-PhD achieved 100% accuracy on Russian news and favored human-written text, while the other two annotators, with only medium accuracy, preferred machine-generated text. For datasets with medium or low detection accuracy (Russian and Arabic summaries, Arabic tweets), machine text was preferred in $\geq\frac{2}{3}$ of the cases. These suggest that when annotators can clearly identify human-written content, social or psychological bias toward peer-authored text emerges, in contrast, when uncertain, their intuitive preferences tend to favor machine outputs.
@@
Are Humans Good Emotional Support?
Not necessary. Interestingly, for emotion-rich questions (QA-emo), where human responses are expected to be more empathetic and be preferred, three out of four annotators actually preferred the machine-generated responses. The remaining annotator disliked both human and machine answers in 22% of the cases. The annotator feedback suggested that this was influenced by the presence of mean-spirited responses from Zhihu users, where some human answers expressed personal biases and lacked empathy (see Appendix G.1).
@@
Human-written essays tend to exhibit greater coherence and sincerity. In contrast, machine-generated essays often lack cohesion, featuring abrupt transitions and sometimes repeating the title in an attempt to simulate continuity. While LLMs can generate related concepts, they struggle to construct a compelling narrative or to seamlessly connect multiple stories under a unified theme. As a result, their outputs often feel shallow—-relying on abstract concepts and rhetorical flourishes without offering substantive insight. The text tends to be verbose yet superficial, giving an impression of grandiosity without meaning. Moreover, LLM-generated essays often adopt an instructive tone, resembling answers to questions rather than peer-level discourse. From a literary perspective, human-written articles often charm with playfulness, sincerity, or novel insights, using precise language to evoke emotion and poetic depth. In contrast, a piece generated by an LLM lacks this literary nuance, leaving little room for contemplation or lingering thoughts. The expression is immediately clear, requiring no further reflection, resembling more of a lecture than an engaging discourse.
@@
Human texts are not always preferred. This inspires us to reflect the ultimate goal of building LLMs that are human-like vs. liked-by-human. The goal of being human-like has a single target, i.e., mimicking human behavior, while to be liked-by-humans involves optimization towards billions of local optima, each shaped by individual preferences. Current language models establish a foundation by learning from human data towards being human-like. As they get more advanced, they can further adapt by incorporating personal data, thus transitioning from merely imitating human behavior to aligning with individual user preferences, and moving from human-like to liked-by-human. Our preference annotations can serve as a valuable resource for guiding models to match individual preferences in multilingual contexts. Also, the data can facilitate investigation of the relationship between annotator characteristics (e.g., MBTI personality, gender, and age) and preferences.
@@ 6 Conclusion and Future Work
We conducted a comprehensive study to investigate the upper bound capability of humans to detect text generated by SOTA LLMs. Based on 16 datasets spanning 9 languages, 9 domains, and 11 LLMs, 19 expert annotators’ accuracy ranged in 50-100%, with an average of 87.6%, revealing that it is not that challenging for experts to identify MGTs. We found that the major gaps between human-written and machine-generated text lie in concreteness, cultural nuances, diversity in length, structure, style, and sentiment, formatting and mixture of languages in the generations. Prompting by explicitly indicating the distinctions could partially bridge this gap, but cultural nuances and diversity remained a challenge. Yet, humans favored machine-generated text in over half of the cases, particularly when they could not recognize which text was written by a human. In future work, we plan to explore human detection accuracy and preference across diverse population groups, especially their relationship with individual characteristics. How do an individual’s philosophy, personality, and experience influence the detection ability and preferences.
@@
Annotator Diversity
All annotators involved in the detection process are advanced NLP researchers, including Msc. and PhD. students, as well as postdocs, specializing in LLMs. No laymen participated in the annotation process, which ensures that the findings and the conclusions were relevant within the expert domain. It is possible that different results could be obtained if laypersons were involved in the detection tasks. However, given our aim to explore the upper bound of human capability in detecting machine-generated text, the current setting is appropriate. Future work could consider involving a broader range of participants to assess the average detection capability across different population groups.
Automatic Linguistic Analysis
We did not include automated linguistic analysis using tools and techniques in this study. The reason lies in that our primary focus is on human-centered case study: how humans perceive quality of the output, identify distinguishable signals between human and machine, detection capabilities and preferences. There have been many studies exploring automatic MGT detection and performing linguistic analysis by automatic methods, while our aim is to understand human perception. Appendix D provides detailed linguistic differences (distinguishable signals) between human-written and machine-generated text which were manually summarized by expert human annotators. They are valuable insights to guide the future LLM improvement. Our future work may explore more differences between human linguistics perceptions and insights derived from advanced computational analysis, across the three text types (human, original machine-generated, and improved machine-generated). Analysis would involve vocabulary features, part-of-speech tagging, dependency structures, sentiment analysis, language model perplexity, and other relevant metrics.
@@
Our study engaged human evaluators in distinguishing between human-written and machine-generated texts, while expressing their preference. All annotators gave informed consent, were fully aware of the study’s objectives, and had the right to withdraw at any time. Since this study was designed to be evaluated by experts, all annotators participating in it were advanced NLP researchers (MSc and PhD students, as well as postdocs, specializing in LLMs). While this ensures a rigorous assessment of the detection capabilities within an expert domain, we acknowledge that findings may not generalize to laypeople. Future research should include a more diverse range of participants to evaluate human detection performance across various demographic and professional backgrounds.
@@
  • Shamardina et al. (2022) Tatiana Shamardina, Vladislav Mikhailov, Daniil Chernianskii, Alena Fenogenova, Marat Saidov, Anastasiya Valeeva, Tatiana Shavrina, Ivan Smurov, Elena Tutubalina, and Ekaterina Artemova. 2022. Findings of the ruatd shared task 2022 on artificial text detection in russian. arXiv preprint arXiv:2206.01583.
  • Shamardina et al. (2025) Tatiana Shamardina, Marat Saidov, Alena Fenogenova, Aleksandr Tumanov, Alina Zemlyakova, Anna Lebedeva, Ekaterina Gryaznova, Tatiana Shavrina, Vladislav Mikhailov, and Ekaterina Artemova. 2025. Coat: Corpus of artificial texts. Natural Language Processing, 31(1):150–175.
@@ A.1 Human Detection on MGT
Early studies probing the outputs of less advanced generative models found that human evaluators could detect the differences. For example, Garbacea et al. (2019) reported that individual evaluators obtained 66.61% accuracy when evaluating online product reviews generated by 12 sequential models before 2019, and majority voting of five annotators’ predictions improved to 72.63%. Ippolito et al. (2020) found that trained evaluators were able to detect GPT2-large text 71.4%. However, since the emergent of GPT-3, it has become challenging for human evaluators to distinguish between human-written and machine-generated text. Evaluators could guess GPT-3-generated stories, news articles and recipes with 50% accuracy (Clark et al., 2021). 203 crowd-soured workers employed from Prolific show 57% average accuracy when discerning news generated by GPT-3.5-turbo and news collected from trusted news organizations, e.g., New York Times, Wall Street Journal, and they obtained 78% on social media comments, both with a wide spread in performance across individuals (Chein et al., 2024). Guo et al. (2023) also showed large differences in detection capability between experts and laymen, where the former is much higher than the later across different domains. Average accuracy of experts is above 90%, but random guess for amateurs, i.e., 48% and 54% for English and Chinese respectively. With expert annotators, academic paper abstracts are harder to detect than the Reddit answers. Detection accuracy ranges from 41 to 94 for Reddit with an average of 77, while 72 for abstracts with individual performance ranging from 60 to 84, revealing that scientific materials are more challenging even for experts who are researchers and frequent users of LLMs Wang et al. (2024b). In addition to binary detection human vs. AI, Wang et al. (2024a) performed 4-class detection to identify which LLM generated a given text, achieving accuracy of 21.2%, below the random guess of 25%, though annotators are all NLP PhD students who are exploring machine-generated text in their research. Dugan et al. (2023) asked annotators to select boundary sentences between human- and machine-authored snippets given a mixed collaborative text by a game tool. They found that players correctly identified the boundary sentence 23.4% of the time (chance being 10%). As shown in Table 1, previous studies primarily focus on English and outputs of GPT-3.5-turbo, non-English languages and more advanced models have been under-investigated. The accuracy of human detection is influenced by languages, generative models, and domains of the given text, as well as the annotators. Large variance is observed among individual evaluators, particularly between experts and laymen. Therefore, we aim to fill this gap by a comprehensive case study exploring whether human can safeguard against AI over 16 dataset spanning nine languages, nine domains and 11 SOTA LLMs in four evaluation setups.
@@ A.3 Detection Accuracy and Preference
What factors may influence individuals’ discerning accuracy? Chein et al. (2024) investigated whether psychological attributes (e.g., fluid intelligence, executive functioning, empathy) or experience (e.g., smartphone and social media habits) predict the ability to differentiate between human and AI materials. Results show that only fluid intelligence — the capacity to reason and solve novel problems without relying on previously acquired knowledge, was strongly associated with the ability to differentiate human from AI materials. Other three factors were not significantly related to the overall detection accuracy. [^1] Fluid intelligence involves the ability to think abstractly, analyze patterns, make decisions in unfamiliar situations, and adapt to new challenges. Fluid intelligence is typically thought to be influenced by genetic and biological factors, and it tends to peak in early adulthood before gradually declining with age. It’s often contrasted with crystallized intelligence, which is based on accumulated knowledge and experience. They also find that increased exposure to mixed AI and human materials online does not enhance one’s ability to differentiate human from AI materials, but rather, may make AI-generated content appear more human-like. Combine together, language proficiency, cognitive intelligence, and familiarity with AI outputs may affect detection accuracy. Intuitively, a high level of language proficiency will make it easier to spot overly formal, repetitive, or mechanically structured phrasing typical of AI outputs. Native speakers with strong awareness of cultural nuance can better understand cultural and idiomatic expressions, facilitating identify content that lacks human-like subtlety or creativity. Strong analytical skills can help detecting inconsistencies, logical flaws, style patterns that indicate MGT. Individuals with frequent interaction with LLMs may recognize their writing patterns like positive tone bias Guo et al. (2023); Chein et al. (2024). Therefore, to probe the upper bound of human capability of detecting current SOTA LLM generations, all annotators for MGT detection and preference labeling are native speakers for each language, Masters’, who are PhD and postdoc students studying LLMs, with 5 female and 14 male.
Приложение B Annotation Tool
To mitigate potential labeling biases arising from raw spreadsheet annotation and to enhance efficiency, we implemented two methods with optimized interfaces and workflows for our annotation: (1) a custom pipeline using the Google Workspace suite, including Apps Script, Google Sheets, and Google Forms. The core idea was to store all data in Google Sheets, use Apps Script to extract data and generate a survey in Google Forms; [^2] https://developers.google.com/apps-script and (2) Label Studio, an open-source multi-type data labeling and annotation tool with a standardized output format. We designed a custom template for our annotation task and collected results using this platform. The annotators were given the choice to use their preferred tool. [^3] https://github.com/HumanSignal/label-studio/
Приложение C Datasets
@@ C.1 Arabic
The SANAD news dataset Einea et al. (2019) is an extensive collection of Arabic news articles in Modern Standard Arabic (MSA), designed to support various Arabic NLP tasks. SANAD dataset comprises over $190,000$ articles from three prominent news websites: AlKhaleej, AlArabiya, and Akhbarona, and it is released under the license of CC-BY-4.0. For this study, we generated a total of $500$ news articles using titles from the SANAD news dataset, with $100$ articles for each of five selected news category topics present in SANAD spanning finance, politics, sports, medicine, and technology. The articles were generated using the OpenAI’s GPT-4o and the prompt used to generate these articles is shown below:
@@ C.2 Chinese
C.3 English Peer Meta Review
@@ C.4 Hindi News
C.5 Italian DICE News
@@ C.9 Vietnamese
Appendix D Distinguishable Signals
@@ D.2 Chinese
D.3 English Peer Meta Review
@@ D.9 Vietnamese
Appendix E Fill the Gap by Prompting?